# IPO: Iterative Preference Optimization for Text-to-Video Generation

### v3

@@ B.3 Impact of Critic Model Size on RL
We conducted a comparative analysis of critic models of different sizes in the iterative preference optimization training process. As shown in Tab. 7, larger critic models provide more precise reward estimations, leading to improved alignment with human preferences. This enhanced reward accuracy contributes to better optimization during preference alignment training, ultimately resulting in superior model performance. Our findings highlight the importance of scaling the critic model to achieve more reliable preference optimization and reinforce the benefits of leveraging larger critic architectures for improved alignment outcomes.
Appendix C Additional Qualitative Comparison
@@
We provide a more detailed qualitative comparison in Fig. 7, which demonstrates the superiority of our model over the baseline. This visualization highlights key improvements in fidelity and overall quality.By comparing outputs side by side, we illustrate the advantages of our model in handling complex scenarios, reinforcing its effectiveness in real-world applications.
Appendix D Societal Impacts
@@ Appendix E Limitations
Figure 7: Enhanced qualitative comparison. We conduct a qualitative comparison between CogVideo-2B and IPO-2B to evaluate their performance

### v4

@@ G.6 Impact of Reward Model Size on RL
We conducted a comparative analysis of reward models of different sizes in the iterative preference optimization training process. As shown in Tab. A3, larger reward models provide more precise reward estimations, leading to improved alignment with human preferences. This enhanced reward accuracy contributes to better optimization during preference alignment training, ultimately resulting in superior model performance. Our findings highlight the importance of scaling the critic model to achieve more reliable preference optimization and reinforce the benefits of leveraging larger critic architectures for improved alignment outcomes.
Appendix H Additional Qualitative Comparison
@@
We provide a more detailed qualitative comparison in Fig. A4 and Fig. A5, which demonstrates the superiority of our model over the baseline. This visualization highlights key improvements in fidelity and overall quality.By comparing outputs side by side, we illustrate the advantages of our model in handling complex scenarios, reinforcing its effectiveness in real-world applications.
Appendix I Use of LLM
Large Language Models (LLMs) were employed solely as auxiliary tools to improve the efficiency of manuscript preparation. Their use included assisting in retrieving related literature, editing and formatting mathematical expressions, polishing the language style, debugging minor code or visualization scripts, and formatting references.
Appendix J Societal Impacts
@@ Appendix I Use of LLM
Figure A4: Enhanced qualitative comparison. We conduct a qualitative comparison between CogVideo-2B and Dual-IPO-2B to evaluate their performance
Figure A5: Enhanced qualitative comparison. We conduct a qualitative comparison between CogVideo-2B and Dual-IPO-2B to evaluate their performance